# Amortized Bayesian Workflow

### v2

@@ Appendix A Background
The ChEES-HMC algorithm [36] is a massively parallel and adaptive extension of Hamiltonian Monte Carlo (HMC) designed to leverage single-instruction multiple-data (SIMD) hardware accelerators such as GPUs. This enables rapid generation of posterior draws following an initial warm-up phase. During warm-up, ChEES-HMC adaptively tunes the trajectory length $T$ and step size $\epsilon$ by maximizing the “Change in the Estimator of the Expected Square” (ChEES), a heuristic that serves as a proxy for reducing autocorrelation in the second moments of the Markov chain. ChEES-HMC is particularly suitable for our amortized workflow, as we can easily generate a large number of good starting points (amortized draws) to launch many short MCMC chains. For our experiments, we used ChEES-HMC to run 2048 parallel chains, organized into 16 superchains with 128 subchains each. This configuration is essential for computing the nested $\smash{\widehat{R}}$ diagnostic [50], which assesses convergence across a large number of short chains.
Appendix B Inference phase: Step 1 diagnostics
In this section, we provide details on the out-of-distribution (OOD) checking procedure used to detect atypical datasets in Step 1 of the workflow. We also discuss potential additional diagnostics that can be used to assess the quality of amortized inference beyond the default check.
B.1 Testing for atypicality in Step 1
Inspired by an out-of-distribution checking method for amortized inference under model misspecification [71], we use a sampling-based hypothesis test to flag atypical datasets where the trustworthiness of amortized inference might be impeded.
Concretely, let $\{y^{(m)}\}_{m=1}^{M}$ denote the training datasets used during simulation-based amortization, and let $h_{\psi}(y)$ denote a low-dimensional summary statistic (e.g., learned or handcrafted) that maps a dataset $y$ into a feature space $\mathbb{R}^{d}$. We compute the summary statistics $s^{(m)}=h_{\psi}(y^{(m)})$ for all training datasets and use them to estimate the empirical mean $\mu_{s}$ and covariance $\Sigma_{s}$ in the summary space. Then, for any dataset $y$, its Mahalanobis distance to the training set is:
$D_{M}(y)=\sqrt{(h_{\psi}(y)-\mu_{s})^{\top}\Sigma_{s}^{-1}(h_{\psi}(y)-\mu_{s})}.$ (8)
To form a null distribution of typical distances, we compute $\{D_{M}(y^{(m)})\}_{m=1}^{M}$ for all training datasets. Given a new observed dataset $y_{\text{obs}}$, we then compute its Mahalanobis distance $D_{M}(y_{\text{obs}})$ and compare it to the empirical distribution of training distances.
We define the out-of-distribution rejection rule as:
$\text{Reject}_{\text{OOD}}(y_{\text{obs}})=\mathbb{I}\left\{D_{M}(y_{\text{obs}})>\text{Quantile}_{\alpha}\left(\{D_{M}(y^{(m)})\}_{m=1}^{M}\right)\right\},$ (9)
where $\alpha$ is typically set to 95% to flag the top 5% most atypical datasets under the training distribution. The type-I error rate ($1-\alpha$) of this test can be set relatively high to obtain a conservative test that will flag many datasets for detailed investigation in further steps of our workflow.
Note.
In the applications of this paper, we perform the above test in the learnt summary space, that is, $h_{\psi}$ is a summary neural network that learns a low-dimensional representation of the data (see Appendix C for details). [^4] Except for Bernoulli GLM problem, the sufficient summary statistics is known without the need for learning.
B.2 Additional diagnostics
In addition to OOD tests for identifying atypical datasets, data-conditional diagnostics can gauge the reliability of amortized posterior inferences for specific datasets. Examples for such data-conditional diagnostics include posterior simulation-based calibration checking (posterior SBC; [83]) and the local classifier two-sample test (L-C2ST; [44]). These diagnostics each offer distinct advantages and limitations.
The choice to apply these additional diagnostics depends on context-specific factors, including the number of observed datasets, the relative computational cost of simulations versus likelihood evaluations, [^5] For example, if likelihood evaluations are cheap, it is often worthwhile to process to Step 2, where Pareto-smoothed importance sampling can offer more informative and powerful diagnostics. and the dimensionality of the observations. Ultimately, whether amortized posterior draws are deemed acceptable hinges on the accuracy requirements of the specific application. By default, we recommend the OOD test for its simplicity, efficiency, and suitability as a first-line diagnostic.
@@ Appendix C Experiment details
For all experiments, we use a coupling-based normalizing flow implemented in BayesFlow [64]. The flow consists of 6 transformation layers, each comprising an invertible normalization, two affine coupling transformations, and a random permutation between elements. Before entering the coupling flow network as conditioning variables, the observed dataset $y$ is encoded into a lower-dimensional summary statistic $h_{\psi}(y)$ via a summary network $h_{\psi}$. This summary network is implemented either as a DeepSet architecture [98] or a SetTransformer [42], depending on the problem setting. Both architectures are designed to handle permutation-invariant data structures. For the Bernoulli GLM, we bypass the summary network and directly use the known 10-dimensional sufficient statistics [48]. The specific choice of summary network for each application is described in the respective problem descriptions below.
@@
For all problems, the neural network is optimized via the AdamW optimizer [47] with weight decay of $10^{-3}$ and a cosine decay learning rate schedule (initial learning rate of $2.5\times 10^{-4}$, a warmup target of $5\times 10^{-4}$, $\alpha=0$) as implemented in Keras [11]. A global gradient clip norm of $1.5$ is applied. Training is performed with a batch size of $512$ for 300 epochs, [^7] For Bernoulli GLM, we only train for 100 epochs. with cosine decay steps set to the product of batch size and epochs. A held-out validation set is used to monitor optimization and select the best-performing model checkpoint.
@@
Following standard practice in Bayesian computation (e.g., PyMC [57], Stan [8]), we transform parameters to an unconstrained space for inference. The amortized neural approximator is trained to estimate parameters in this unconstrained space. PSIS operates independently of the parameterization and thus remains unaffected by this transformation. ChEES-HMC also performs inference in the unconstrained space. All evaluation metrics (W1 and MMTV distances) are computed in this space. However, parameter recovery and simulation-based calibration plots are shown in the original constrained space for better interpretability.
@@
For all applications, the full workflow—including amortized training, inference, diagnostics, Pareto-smoothed importance sampling, and ChEES-HMC sampling—was conducted on a single NVIDIA V100 GPU (32GB), 8 cores of an AMD EPYC 7452 processor, and 8-16GB RAM. For runtime details across experiments, refer to Table 1 in the main text. The core code base was built using BayesFlow [64] (MIT license), PyMC [57] (Apache-2.0 license), ArviZ [39] (Apache-2.0 license) and JAX [4] (Apache-2.0 license). We used the implementation of ChEES-HMC provided by TensorFlow Probability [16] (Apache-2.0 license).
@@ C.1 Generalized extreme value distribution
Following Caprani [7], the prior distribution for the parameters of the generalized extreme value distribution (GEV) is defined as:
$\mu$ $\sim\mathcal{N}(3.8,0.04)$ (10)
@@
We use $10,000$ simulated parameter–observation pairs for training the amortized approximator, $1000$ for validation, and $200$ for training-phase diagnostics, including parameter recovery and simulation-based calibration.
@@
The closed-world diagnostics (parameter recovery and simulation-based calibration checking) in Figure 7 indicate that the neural network training has successfully converged to an acceptable posterior approximator within the scope of the training set.
@@
In order to emulate distribution shifts that arise in real-world applications while preserving the controlled experimental environment, we simulate the “observed” datasets from a joint model whose prior is 2×2$\times$2 × wider (i.e., with 4×4$\times$4 × the variance) than the model used during training. More specifically, the prior is specified as:
$\mu$ $\sim\mathcal{N}(3.8,0.16)$ (11)
@@ C.2 Bernoulli GLM
Following Lueckmann et al. [48], we set the prior for $\theta$ as:
$\theta\sim\mathcal{N}\left(0,\begin{bmatrix}2&0\\ 0&(F^{\top}F)^{-1}\end{bmatrix}\right),$ (12)
where the matrix $F$ is defined such that $F_{i,i-2}=1$, $F_{i,i-1}=-2$, $F_{i,i}=1+\sqrt{\frac{i-1}{9}}$, and $F_{i,j}=0$ otherwise, for $1\leq i,j\leq 9$. The task duration is set to $T=100$, with fixed input vectors $\{v_{i}\}_{i=1}^{100}$, where each $v_{i}\in\mathbb{R}^{10}$. Corresponding observations are denoted by $\{y_{i}\}_{i=1}^{100}$. Further details can be found in [48, 31].
@@
We use $10,000$ simulated parameter–observation pairs for training the amortized approximator, $1000$ for validation, and $200$ for training-phase diagnostics, including parameter recovery and simulation-based calibration.
@@
The closed-world diagnostics (parameter recovery and simulation-based calibration checking) in Figure 8 indicate that the neural network training has successfully converged to an acceptable posterior approximator within the scope of the training set.
@@
We generate $K=$10,000$$ in-distribution test datasets by sampling parameters from the model prior and simulating corresponding observations $\{y_{i}\}_{i=1}^{100}$ from the Bernoulli distribution.
@@ C.3 Psychometric curve fitting
We adopt an overdispersed psychometric model [75] with the error function (erf) as the sigmoid function in the psychometric function:
@@
[…]
@@
We use $50,000$ simulated parameter–observation pairs for training the amortized approximator, $1000$ for validation, and $200$ for training-phase diagnostics, including parameter recovery and simulation-based calibration.
@@
The closed-world diagnostics (parameter recovery and simulation-based calibration checking) in Figure 9 indicate that the neural network training has successfully converged to an acceptable posterior approximator within the scope of the training set.
@@
Our empirical evaluation uses 8,526 mouse behavioral datasets from the International Brain Laboratory public database [85]. We retrieve the data using the provided API with the argument task="biasedChoiceWorld", which corresponds to behavioral data collected after the mice have completed training. Each dataset is processed into an observation tensor of shape $(9,3)$, where each row contains the number of correct trials $y_{i}$, the total number of trials $n_{i}$, and the stimulus level $x_{i}$.

### v3

@@ Appendix A Background
The ChEES-HMC algorithm (Hoffman et al., 2021) is a massively parallel and adaptive extension of Hamiltonian Monte Carlo (HMC) designed to leverage single-instruction multiple-data (SIMD) hardware accelerators such as GPUs. This enables rapid generation of posterior draws following an initial warm-up phase. During warm-up, ChEES-HMC adaptively tunes the trajectory length $T$ and step size $\epsilon$ by maximizing the “Change in the Estimator of the Expected Square” (ChEES), a heuristic that serves as a proxy for reducing autocorrelation in the second moments of the Markov chain. ChEES-HMC is particularly suitable for our amortized workflow, as we can easily generate a large number of good starting points (amortized draws) to launch many short MCMC chains. For our experiments, we used ChEES-HMC to run 2048 parallel chains, organized into 16 superchains with 128 subchains each. This configuration is essential for computing the nested $\smash{\widehat{R}}$ diagnostic (Margossian et al., 2024), which assesses convergence across a large number of short chains. The pseudocode explaining the use of ChEES-HMC in Step 3 is provided in Algorithm 6.
Algorithm 6 Use of ChEES-HMC in Step 3
1:Log-posterior density $\log p(\theta,y_{\text{obs}})=\log p(y_{\mathrm{obs}}\,|\,\theta)+\log p(\theta)$ and its gradient w.r.t. $\theta$; number of superchains $K$; number of subchains per superchain $M$; warm-up length $N_{\mathrm{warmup}}$ (e.g., $N_{\mathrm{warmup}}=200$); after warm-up sampling length NN (e.g., $N=1$)
2:Collect $K$ unique amortized posterior draws or PSIS-corrected draws for chain initialization; each of these $K$ draws must have a finite log-posterior density value.
3:For each superchain group $k=1,\dots,K$, initialize $M$ subchains at the same initial state (total $K\times M$ chains)
4:Run ChEES-HMC warm-up for $N_{\mathrm{warmup}}$ iterations to adapt step size $\epsilon$ and trajectory length $L$
5:Fix $(\epsilon,L)$ and run NN iterations to collect draws from KMKM parallel chains
6:Compute nested $\widehat{R}$ for each parameter component of $\theta$
7:If nested $\widehat{R}$ is below a chosen threshold (e.g. <1.01<1.01), accept the combined draws as Step 3 posterior draws; otherwise increase warmup length, use alternative MCMC algorithm (e.g., NUTS-HMC) or revise the model
Nested $\widehat{R}$ diagnostic.
The potential scale reduction factor $\widehat{R}$ (Gelman & Rubin, 1992; Vehtari et al., 2021) is arguably the most popular diagnostic for assessing the convergence of MCMC chains. The basic idea is that multiple MCMC chains starting from overdispersed initial points should produce similar Monte Carlo estimators if they have converged, i.e., the impact of initialization vanishes as the chains converge to the stationary distribution. Nested $\widehat{R}$ diagnostic (Margossian et al., 2024) extends the classical $\widehat{R}$ diagnostic for monitoring convergence of many-short-chain MCMC samplers such as the ChEES-HMC algorithm. It requires organizing chains into $K$ superchains, each consisting of $M$ subchains that share the same initial point. Thus, one can assess convergence through comparing the variability between superchains and the variability within superchains, similar to the standard $\widehat{R}$ diagnostic. We provide the pseudocode for computing the nested $\widehat{R}$ diagnostic in Algorithm 7 for reference.
Algorithm 7 Nested $\widehat{R}$ diagnostic
1:Posterior draws $\{\theta^{(nmk)}\}$ after warm-up, where $k\in\{1,\dots,K\}$ (superchains), $m\in\{1,\dots,M\}$ (subchains), $n\in\{1,\dots,N\}$ (draws); scalar function of interest $f$
2:Compute scalar values $f^{(nmk)}\leftarrow f(\theta^{(nmk)})$ for all $n,m,k$.
3:Compute subchain means: $\bar{f}^{(\cdot mk)}\leftarrow\frac{1}{N}\sum_{n=1}^{N}f^{(nmk)}$
4:Compute superchain means $\bar{f}^{(\cdot\cdot k)}\leftarrow\frac{1}{M}\sum_{m=1}^{M}\bar{f}^{(\cdot mk)}$ and overall mean $\bar{f}^{(\cdot\cdot\cdot)}\leftarrow\frac{1}{K}\sum_{k=1}^{K}\bar{f}^{(\cdot\cdot k)}$
5:for each superchain $k=1,\dots,K$ do
6:  Compute between-chain variance $\tilde{B}_{k}$:
7:   If $M>1$, $\tilde{B}_{k}\leftarrow\frac{1}{M-1}\sum_{m=1}^{M}(\bar{f}^{(\cdot mk)}-\bar{f}^{(\cdot\cdot k)})^{2}$; else $\tilde{B}_{k}\leftarrow 0$.
8:  Compute within-chain variance $\tilde{W}_{k}$:
9:   If $N>1$, $\tilde{W}_{k}\leftarrow\frac{1}{M}\sum_{m=1}^{M}\left(\frac{1}{N-1}\sum_{n=1}^{N}(f^{(nmk)}-\bar{f}^{(\cdot mk)})^{2}\right)$; else $\tilde{W}_{k}\leftarrow 0$.
10:end for
11:Compute between-superchain variance: $\widehat{B}_{\nu}\leftarrow\frac{1}{K-1}\sum_{k=1}^{K}(\bar{f}^{(\cdot\cdot k)}-\bar{f}^{(\cdot\cdot\cdot)})^{2}$
12:Compute within-superchain variance: $\widehat{W}_{\nu}\leftarrow\frac{1}{K}\sum_{k=1}^{K}(\tilde{B}_{k}+\tilde{W}_{k})$
13:Return $\text{nested }\widehat{R}\leftarrow\sqrt{\frac{\widehat{W}_{\nu}+\widehat{B}_{\nu}}{\widehat{W}_{\nu}}}$
Appendix B Best practices for training amortized estimators
Amortized inference approaches problems of Bayesian modeling with methods from deep learning. While the precise training setup naturally depends on the concrete problem at hand, some general principles have proven help across a wide range of amortized inference applications. We summarize these here as initial guidance for applied practitioners.
Rely on established tooling.
Modern libraries for amortized inference such as sbi (Boelts et al., 2025) and BayesFlow (Radev et al., 2023) provide well-tested neural density estimators, training loops, and data pipelines. In many cases, their default architectures and optimization settings already yield strong performance without manual tuning. First and foremost, we strongly recommend starting from these defaults and only introducing additional complexity if the diagnostics indicate deficiencies.
Monitor the training process.
In many amortized inference settings, data are generated on the fly by a forward simulation program (see Eq. 2), and training does not rely on a fixed dataset. In this case, classical data splits into training and validation set are less meaningful, since each minibatch effectively constitutes fresh data from the joint model. Nonetheless, it is still important to monitor training progress with multiple signals, such as the training loss, calibration diagnostics, or summary statistics of posterior samples. These checks help assess whether the model continues to improve or has already reached a performance plateau after a short period. In other settings, amortized inference may be trained on a fixed set of simulated data, for example due to an expensive simulator or precomputed datasets. In such cases, holding out a validation set is strongly recommended to detect overfitting and guide selection of the amortized estimator.
Track multiple performance signals.
Regardless of whether data are simulated on the fly or fixed, we recommend monitoring multiple indicators during training (e.g., after each epoch). Loss curves provide a coarse signal of optimization progress but are often insufficient on their own. For example, normalizing flows are usually trained with a negative log-likelihood loss, which does not account for mode coverage in multi-modal posterior distributions. Complementary diagnostics such as simulation-based calibration, parameter recovery, or posterior predictive checks on held-out datasets offer more direct insight into posterior quality and failure modes.
Assess and adjust model expressiveness.
When training stagnates or diagnostics indicate systematic errors, the limiting factor is often model expressiveness rather than optimization details. Underexpressive models may show symptoms such as poor parameter recovery, persistent miscalibration, or posterior collapse toward the prior (i.e., data insensitivity), even when the training loss decreases. A pragmatic strategy is to begin with an overly expressive architecture (i.e., many trainable weights) to establish a performance baseline, and then gradually simplify the model. Conversely, if diagnostics remain unsatisfactory, increasing model capacity is often more effective than tuning hyperparameters of the optimizer.
Accept that training is iterative.
Even with modern tooling, amortized inference training may require many iterations during development, especially for complex or weakly identifiable models. The objective is not to find a universally optimal neural architecture, but to reach a regime where the amortized posterior is reliable enough to enter the adaptive workflow proposed in this paper, where subsequent diagnostics and correction steps can take over.
@@ Appendix C Experiment details
For all experiments, we use a coupling-based normalizing flow implemented in BayesFlow (Radev et al., 2023). The flow consists of 6 transformation layers, each comprising an invertible normalization, two affine coupling transformations, and a random permutation between elements. Before entering the coupling flow network as conditioning variables, the observed dataset $y$ is encoded into a lower-dimensional summary statistic $h_{\psi}(y)$ via a summary network $h_{\psi}$. This summary network is implemented either as a DeepSet architecture (Zaheer et al., 2017) or a SetTransformer (Lee et al., 2019), depending on the problem setting. Both architectures are designed to handle permutation-invariant data structures. For the Bernoulli GLM, we bypass the summary network and directly use the known 10-dimensional sufficient statistics (Lueckmann et al., 2021). The specific choice of summary network for each application is described in the respective problem descriptions below.
@@
For all problems, the neural network is optimized via the AdamW optimizer (Loshchilov & Hutter, 2019) with weight decay of $10^{-3}$ and a cosine decay learning rate schedule (initial learning rate of $2.5\times 10^{-4}$, a warmup target of $5\times 10^{-4}$, $\alpha=0$) as implemented in Keras (Chollet et al., 2015). A global gradient clip norm of $1.5$ is applied. Training is performed with a batch size of $512$ for 300 epochs, [^9] For Bernoulli GLM, we only train for 100 epochs. with cosine decay steps set to the product of batch size and epochs. A held-out validation set is used to monitor optimization and select the best-performing model checkpoint.
@@
Following standard practice in Bayesian computation (e.g., PyMC; Oriol et al., 2023, Stan; Carpenter et al., 2017), we transform parameters to an unconstrained space for inference. The amortized neural estimator is trained to estimate parameters in this unconstrained space. PSIS operates independently of the parameterization and thus remains unaffected by this transformation. ChEES-HMC also performs inference in the unconstrained space. All evaluation metrics (W1 and MMTV distances) are computed in this space. However, parameter recovery and simulation-based calibration plots are shown in the original constrained space for better interpretability.
@@
For all applications, the full workflow—including amortized training, inference, diagnostics, Pareto-smoothed importance sampling, and ChEES-HMC sampling—was conducted on a single NVIDIA V100 GPU (32GB), 8 cores of an AMD EPYC 7452 processor, and 8-16GB RAM. For runtime details across experiments, refer to Table 1 in the main text. The core code base was built using BayesFlow (Radev et al., 2023) (MIT license), PyMC (Oriol et al., 2023) (Apache-2.0 license), ArviZ (Kumar et al., 2019) (Apache-2.0 license) and JAX (Bradbury et al., 2018) (Apache-2.0 license). We used the implementation of ChEES-HMC provided by TensorFlow Probability (Dillon et al., 2017) (Apache-2.0 license).
@@ C.1 Generalized extreme value distribution
Following Caprani (2021), the prior distribution for the parameters of the generalized extreme value distribution (GEV) is defined as:
[…]
@@
We use $10,000$ simulated parameter–observation pairs for training the amortized estimator, $1000$ for validation, and $200$ for training-phase diagnostics, including parameter recovery and simulation-based calibration.
@@
The closed-world diagnostics (parameter recovery and simulation-based calibration checking) in Figure 7 indicate that the neural network training has successfully converged to an acceptable posterior estimator within the scope of the training set.
@@
In order to emulate distribution shifts that arise in real-world applications while preserving the controlled experimental environment, we simulate the “observed” datasets from a joint model whose prior is $2\times$ wider (i.e., with $4\times$ the variance) than the model used during training. More specifically, the prior is specified as:
[…]
@@ C.2 Bernoulli GLM
Following Lueckmann et al. (2021), we set the prior for $\theta$ as:
[…]
where the matrix $F$ is defined such that $F_{i,i-2}=1$, $F_{i,i-1}=-2$, $F_{i,i}=1+\sqrt{\frac{i-1}{9}}$, and $F_{i,j}=0$ otherwise, for $1\leq i,j\leq 9$. The task duration is set to $T=100$, with fixed input vectors $\{v_{i}\}_{i=1}^{100}$, where each $v_{i}\in\mathbb{R}^{10}$. Corresponding observations are denoted by $\{y_{i}\}_{i=1}^{100}$. Further details can be found in Lueckmann et al. (2021); Gonçalves et al. (2020).
@@
We use $10,000$ simulated parameter–observation pairs for training the amortized estimator, $1000$ for validation, and $200$ for training-phase diagnostics, including parameter recovery and simulation-based calibration.
@@
The closed-world diagnostics (parameter recovery and simulation-based calibration checking) in Figure 8 indicate that the neural network training has successfully converged to an acceptable posterior estimator within the scope of the training set.
@@
We generate K=10, [^0] K=\$10,000\$ in-distribution test datasets by sampling parameters from the model prior and simulating corresponding observations $\{y_{i}\}_{i=1}^{100}$ from the Bernoulli distribution.
@@ C.3 Psychometric curve fitting
We adopt an overdispersed psychometric model (Schütt et al., 2016) with the error function (erf) as the sigmoid function in the psychometric function:
@@
where $n_{i}$ denotes the number of trials, and $x_{i}$ is the stimulus level. Stimuli are presented at nine fixed levels: $x_{i}\in\{-100.0,\ -25.0,\ -12.5,\ -6.25,\ 0.0,\ 6.25,\ 12.5,\ 25.0,\ 100.0\}$ and each value is further normalized by dividing by 100.
@@
We use $50,000$ simulated parameter–observation pairs for training the amortized estimator, $1000$ for validation, and $200$ for training-phase diagnostics, including parameter recovery and simulation-based calibration.
@@
The closed-world diagnostics (parameter recovery and simulation-based calibration checking) in Figure 9 indicate that the neural network training has successfully converged to an acceptable posterior estimator within the scope of the training set.
@@
Our empirical evaluation uses 8,526 mouse behavioral datasets from the International Brain Laboratory public database (The International Brain Laboratory et al., 2021). We retrieve the data using the provided API with the argument task="biasedChoiceWorld", which corresponds to behavioral data collected after the mice have completed training. Each dataset is processed into an observation tensor of shape $(9,3)$, where each row contains the number of correct trials $y_{i}$, the total number of trials $n_{i}$, and the stimulus level $x_{i}$.